OpenStudy (anonymous):

@hero Factor the algebraic expression below in terms of a single trigonometric function. sin 2x + sin x - 2

OpenStudy (marissalovescats):

Sin^2x+sinx-2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Do you know your trig identities @Nairz ????

OpenStudy (anonymous):

Yes

OpenStudy (marissalovescats):

Okay so it is the same if you had x^2+x-2 That would be (x+2)(x-1) right? So (sinx+2)(sinx-1) Make sense?

OpenStudy (anonymous):

\[\Large \sin 2x + \sin x - 2~or~\sin^2x+sinx-2?\]

OpenStudy (marissalovescats):

With problems like this, it helps if I just look at it in terms of x

OpenStudy (anonymous):

Ok, forget my help lol thanks @marissalovescats

OpenStudy (anonymous):

Thanks, guys and gals!

OpenStudy (marissalovescats):

You're welcome! :)

OpenStudy (anonymous):

SO the answer is the factored version? @marissalovescats

OpenStudy (marissalovescats):

Yes

OpenStudy (mathmale):

Caution, @Nairz: You were fortunately that Marissa was able to interpret your presentation of this problem correctly. Your sin 2x + sin x - 2 actually means "the sine of twice the angle x plus the sine of x minus 2," when in fact you meant "the square of the sine of x plus the sine of x minus 2. Please use parentheses to be clear: (sin x)^2, or \[\sin ^{2}x,~or~(\sin x)^2.\]

OpenStudy (mathmale):

Math is an exact science, so the more clearly you can communicate (or read) it, the better your chances of doing well in this subject.